0.05:X as 0.15:30

X=

Answer by stanbon(75887)

0.05:X as 0.15:30
-----
Equation:
0.05/x = 0.15/30
---
x = 0.05*30/0.15
---
x = 10
=============
Cheers,
Stan H.